James Dean trades in clothing for the middle market. James Dean Clothing use the Periodic Inventory Method to keep records of trade goods. The following is an extract of balances from the general ledger for the year ended 28 February 2021

extract of balances as at28 february 2021

R

Revenue from sale of trade goods

735 100

Rent income received from tenant

66 000

Inventory: Trade goods - balance at 1 March 2020

: Consumables on hand at 1 March 2020

150 500

7 600

Purchases

353 000

Settlement discounts received

18 400

Settlement discounts granted

15 800

Railage inwards

36 200

Railage outwards

39 100

Goods returned to suppliers

21 400

Goods returned by customers

34 600

Additional information:

1. Physical stock-take at 28 February 2021 determined the following: - Trade goods R166 800 - Consumables on hand R 7 940

2. Goods to the value R7 600 were donated to the local community organization

3. An invoice for R4 500 was received on 28 February 2021 for goods transported from the supplier to James Dean Clothing. The invoice needs to be recorded.

4. James Dean rents office space to a tenant. Rental income of R18 700 was credited to the sales in error. A correction needs to be recorded.

Required:

3.1 Show journal entries to record adjustments for all the additional information. (6)

3.2 Prepare the trading section of the Statement of Profit or Loss and other comprehensive income for James Dean Clothing for the year ended 28 February 2021. Sales, Cost of sales and gross profit must be shown. Also show the gross profit percentage (%) on cost. (16)

3.3 Briefly describe three important features of the Perpetual Inventory Method. (3)



1
Expert's answer
2021-10-22T13:22:46-0400

Solution:

The calculations for 3.1 and 3.2 are as below:





 

Gross profit % on cost = "\frac{GP}{COGS}\times 100 = \frac{166,940}{514,860} \times 100 = 32.42\%"


3.3.). The three important features of Perpetual Inventory Method include the following:

1.). It offers a reliable technique.

2.). It offers less cost and time.

3.). It offers regular and continuous stock checks.